Vera estimated that it will take her 16 2/3 h to complete a project. She spent 4 1/6 h working last weekend. What fraction of the time needed to complete the project did she work last weekend?
Answer by rfer(16322)

(4 1/6)/(16 2/3)=
(25/6)/(50/3)=
25/6*3/50=75/300=1/4